OpenStudy (anonymous):

F(x) = x 2 + 3x - 2 F(a) =

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle F(x) = x^2 + 3x - 2 }\) to find \(\large\color{black}{ \displaystyle F(a) }\) plug in a instead of x.

OpenStudy (solomonzelman):

I mean "a" is still a variable, so that doesn't make it any different.

OpenStudy (solomonzelman):

can you plug in "a" for "x" ?

OpenStudy (solomonzelman):

or want an example?

OpenStudy (anonymous):

f(a)=x^2+3x-2?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

you aren't just replacing the "x" with an "a" in `F(a)`. RATHER, you replace "x" with an "x" EVERYWHERE.

OpenStudy (solomonzelman):

for example, if I wanted to find F(2) \(\large\color{black}{ \displaystyle F(2) = (2)^2 + 3(2) - 2 }\) then I will need to simplify (if that is a possibility) \(\large\color{black}{ \displaystyle F(2) = 4 +6 - 2 }\) \(\large\color{black}{ \displaystyle F(2) = 8 }\)

OpenStudy (anonymous):

I think I get it. At least a little better.

OpenStudy (solomonzelman):

yes, so how (do you think) you will find F(a)

OpenStudy (solomonzelman):

(again, plug in "a" instead of "x" everywhere)

OpenStudy (anonymous):

Order of operations?

OpenStudy (solomonzelman):

no, just substitute "a" for "x"

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle F(x) = x^2 + 3x - 2 }\) \(\large\color{black}{ \displaystyle F(a) =~~.... }\) (what will it equal ?)

OpenStudy (anonymous):

f(a)=a^2=3a-2?

OpenStudy (solomonzelman):

yes, if that second = sign was supposed to be a +

OpenStudy (anonymous):

Thank you =)

OpenStudy (solomonzelman):

you welcome !